OpenStudy (anonymous):

For f(x)=1/x-5 and g(x)=x^2+2 Find the expression for g(x). Substitute the value of g(x) into the function f(x) in place of x to find the value of f(g(x))

OpenStudy (astrophysics):

So we're looking for f(g(x))?

OpenStudy (anonymous):

Looking for g(x) too! But not sure if it's just g(x)=x^2+2 orrr

OpenStudy (astrophysics):

That just means plug in function g(x) where ever there is an x in f(x)

OpenStudy (astrophysics):

g(x) is just x^2+2

OpenStudy (astrophysics):

If your f(x) = 1/(x-5) or is it (1/x)-5

OpenStudy (anonymous):

it's 1/(x-5)

OpenStudy (astrophysics):

Ok, always put brackets! :) So go ahead and find f(g(x)) as I told you how to

OpenStudy (anonymous):

I got 1/(x^2-3)

OpenStudy (astrophysics):

\[f(g(x)) = \frac{ 1 }{ (x^2+2)-5 }\]

OpenStudy (anonymous):

It also has a part 2 that says: (gof)(6) a. find f(6) and b Substitute the value you found in Part 1 into g(x) to find g(f(6))

OpenStudy (astrophysics):

So yes, you're right :)

OpenStudy (anonymous):

thanks!

OpenStudy (astrophysics):

(g o f)(x) is the same thing as g(f(x))

OpenStudy (anonymous):

oh, so do i just plug in 6 to 1/(x^2-3) ?

OpenStudy (astrophysics):

So plug in function f(x) in g(x) then plug in 6 for (g o f)(6)

OpenStudy (astrophysics):

No, that's f(g(x))

OpenStudy (anonymous):

oh so is it 1/(x-5) + 2

OpenStudy (astrophysics):

\[g(f(x)) = \left( \frac{ 1 }{ x-5 } \right)^2+2\]

OpenStudy (anonymous):

Do i need to simplify that? or no

OpenStudy (astrophysics):

Just find g(f(6))

OpenStudy (anonymous):

wait, is Find F(6) just plugging in 6 to f(x)

OpenStudy (anonymous):

and g(f(6)) is the equation you gave? for part b

OpenStudy (astrophysics):

\[g(f(x)) = \left( \frac{ 1 }{ x-5 } \right)^2+2\] \[g(f(6)) = \left( \frac{ 1 }{ 6-5 } \right)^2+2\]

OpenStudy (anonymous):

So it's 3 for the question that asks: Substitute the value you found in Part 1 into g(x) to find g(f(6))

OpenStudy (anonymous):

and there's another question that says find f(6) so would that just be 1

OpenStudy (astrophysics):

Can you post the question, it's a bit confusing with all the notation

OpenStudy (astrophysics):

Especially when you're not using LaTeX

OpenStudy (anonymous):

1) Find f(6). 2) Substitute the value you found in Part 1 into g(x) to find g(f(6))

OpenStudy (astrophysics):

I mean take an image of the question

OpenStudy (astrophysics):

and post it here

OpenStudy (anonymous):

oh sorry! hold on

OpenStudy (astrophysics):

Yeah, I'm not sure which question is connected to what, so it's a bit confusing :P

OpenStudy (anonymous):

it was separated into two pages, sorry!

OpenStudy (astrophysics):

Well I'm not sure why you didn't just take a picture of the full page, but it seems incomplete, your question for part A seems as if it wants you to find a expression using g(x) from the graph.

OpenStudy (anonymous):

Oh, sorry ignore the graph, it's a different question. sorry!!

OpenStudy (astrophysics):

Huh? Then this really makes no sense, are the pages both completely different questions?

OpenStudy (anonymous):

Nope, they're supposed to go together

OpenStudy (astrophysics):

Oh I see, the graph is on a different piece of paper!

OpenStudy (anonymous):

yeah!! sorry haha :/

OpenStudy (astrophysics):

Ok, so lets do it all over again

OpenStudy (astrophysics):

We're given \[f(x) = \frac{ 1 }{ x-5 } ~~~\text{and}~~~~g(x) = x^2+2 \] Part 1, A seems they just want you to find the expression for g(x) meaning they are just seeing if you understand the question, so it's just \[g(x) = x^2+2\] part B wants you to find the f(g(x)) so we take function g(x) and plug it into f(x) \[f(g(x)) = \frac{ 1 }{ (x^2+2)-5 }\]

OpenStudy (astrophysics):

You can do the simplifications, now lets move on to part 2

OpenStudy (astrophysics):

We need to find f(6) that just means we need to find \[f(6) = \frac{ 1 }{ 6-5 }\] which gives us what?

OpenStudy (anonymous):

1!

OpenStudy (astrophysics):

Good

OpenStudy (anonymous):

and so B would be 3 right?

OpenStudy (astrophysics):

Lets see

OpenStudy (astrophysics):

It's asking us to substitute the value we found in part 1, into g(x) so we can find g(f(6))

OpenStudy (anonymous):

yup! so I would just find g(f(x)) right?

OpenStudy (astrophysics):

What we found in part 1 was \[g(x) = x^2+2\]

OpenStudy (astrophysics):

Is the one they are referring to I believe

OpenStudy (astrophysics):

So all you need to do here is, find g(f(x)) first then g(f(6))

OpenStudy (astrophysics):

\[g(f(x)) = \left( \frac{ 1 }{ x-5 } \right)^2+2\]

OpenStudy (anonymous):

I got 3

OpenStudy (astrophysics):

\[g(f(6)) = \left( \frac{ 1 }{ 6-5 } \right)^2+2\]

OpenStudy (anonymous):

so yes, 3?

OpenStudy (astrophysics):

Yeah

OpenStudy (anonymous):

Thanks so much for your help!! :))

OpenStudy (astrophysics):

So what this question is trying to get across is, you knowing what the notation means and what exactly these compositional functions are doing. So notice we actually took what we solved f(6) and just plugged in g(x)

OpenStudy (anonymous):

I get it now haha :) thanks!!

OpenStudy (astrophysics):

We could've very well put \[g(f(6)) = f(6)^2 + 2 = 1^2+2\]

OpenStudy (astrophysics):

Np